OpenStudy (anonymous):

16^(5/4) Help with steps Please and thanks

OpenStudy (zzr0ck3r):

\(16^{\frac{5}{4}}=(\sqrt[4]{16})^5\) What is \(\sqrt[4]{16}\) ?